The complete combustion of 1 mole of propane (C3H8) is represented by the reaction shown. C3H8 + 5O2 → 3CO+ 4H2O How many grams of water are produced in this reaction?


1
Expert's answer
2022-01-30T08:00:34-0500

From the balanced equation, 4 moles of water was produced in the reaction.

Number of moles = mass/molar mass

Molar mass of water = (1*2)+16

= 18

4 = mass/18

Mass = 4*18

Therefore, the mass of water produced in the reaction is 72g.